I need to know which one of these equations has the same solution as the equation 2x+6=32.
2x=38
x-3=6
x+6=16
2(x-3)=16
2(x+3)=32
I know there is a difference between sloveing, solutions, and simplifying but I just don't know wtah it is?

Answer by mananth(16946)

2x+6=32.
/2
x+3=16

Options
2(x+3)=32 is right
/2
x+3=16